OpenStudy (anonymous):

Determine the gravitational force that you exert on another person 1.50 m away. Assume that you and the other person are point masses of 74.0 kg each.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Newton's_law_of_universal_gravitation F = G(m1 * m2)/r² where G is the gravitational constant Plugging everything in. F = G(74kg * 74kg)/(1.5m)² F = G(74kg * 74kg)/(1.5m)² F = G(5476kg²)/2.25m² F = G(2433.78 kg²/m²) F = 1.62406139 × 10^-7 newtons